OpenStudy (anonymous):

Tristan says that solutions to radical equations are extraneous. Hadley says that solutions to radical equations are non-extraneous. Find an equation where Tristan is correct and a second equation where Hadley is correct. Using complete sentences, explain each step when solving to justify your examples.

OpenStudy (anonymous):

Tristan: √x-2=−3 x−2=9 x=11 Check: √11-2=−3 9√=−3 3≠−3,extraneous

OpenStudy (anonymous):

Hadley: √x-2=3 x−2=9 x=11 Check: √11-2=3 9√=3 3=3,non−extraneous

OpenStudy (anonymous):

does that make sense

OpenStudy (anonymous):

@LexiLuvv2431

OpenStudy (anonymous):

omgg thank youuu soo much. yes it doess @israa88

OpenStudy (anonymous):

Lol no problem :)

OpenStudy (anonymous):

Do You think you could help with another!!??? @israa88

OpenStudy (anonymous):

ill try :)

OpenStudy (anonymous):

@LexiLuvv2431

OpenStudy (anonymous):

Sorryy I fell asleep @israa88 But the question is Debbie and trisha are looking at the equation ' The square root of 4x-6 = the square root of x ' Debbie says that the solution is extraneous. Trisha says the solution is non-extraneous. Is debbie correct? Is Trisha correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.